### v1

# Dual Length Codes for Lossless Compression of BFloat16

Abstract
Training and serving Large Language Models (LLMs) relies heavily on parallelization and collective operations, which are frequently bottlenecked by network bandwidth. Lossless compression using e.g., Huffman codes can alleviate the issue, however, Huffman codes suffer from slow, bit-sequential decoding and high hardware complexity due to deep tree traversals. Universal codes e.g., Exponential-Golomb codes are faster to decode but do not exploit the symbol frequency distributions. To address these limitations, this paper introduces Dual Length Codes, a hybrid approach designed to balance compression efficiency with decoding speed. Analyzing BFloat16 tensors from the Gemma model, we observed that the top 8 most frequent symbols account for approximately 50% of the cumulative probability. These 8 symbols are assigned a short 4 bit code. The remaining 248 symbols are assigned a longer 9 bit code. The coding scheme uses a single prefix bit to distinguish between the two code lengths. The scheme uses a small Look Up Table with only 8 entries for encoding and decoding. The scheme achieves a compressibility of 18.6% in comparison to 21.3% achieved by Huffman codes, but it significantly speeds up the decoding and simplifies the hardware complexity.
1 Background
Training and serving Large Language Models (LLMs) e.g., Gemini [3], Gemma [16, 20], LLaMA [22], GPT [18] require partitioning (sharding) the data (parameters, activations, optimizer state etc.) and parallelizing the computation across multiple accelerators. There are multiple paradigms of parallelism e.g., Data Parallelism, Tensor Parallelism, Pipeline Parallelism, Expert Parallelism and Sequence Parallelism [21, 14, 19, 17, 24, 15, 4]. Different parallelization strategies invoke different collective operations e.g., AllReduce, ReduceScatter, AllGather, AlltoAll [5]. Collective operations are typically bounded by network bandwidth. Lossless compression is an effective way to reduce the network traffic and improve collective performance.
Huffman codes [13] either directly or as part of other algorithms e.g., DEFLATE [7], Zstandard [6], Brotli [2] are commonly used for lossless data compression. They exploit the distribution of symbol frequencies and are optimal entropy codes. However, they are variable length and decoding them requires a binary tree traversal. The decoding is slow and bit sequential i.e., the decode latency is proportional to the number of bits in the encoded input.
Universal codes [23] e.g., Elias Gamma codes [9], Elias Delta codes [8], Elias Omega codes [10], Exponential Golomb codes [12] are also variable length codes. In these codes, the length of the code is part of the code, for example, Elias Gamma codes use a series of leading zeros to indicate the length of the code. Therefore, the decoding is not completely bit sequential and hence faster. However, these codes do not exploit the distribution of symbol frequencies and hence are not optimal.
2 Objective
Our objective is to develop codes such that
  • They exploit the distribution of symbol frequencies, similar to Huffman codes.
  • The decoder is not bit sequential and hence faster, similar to Universal codes.
3 Experimental Setup
We analyzed the Gemma 2B model [16] during Supervised Fine Tuning (SFT). The model has 18 layers and is sharded over 64 TPUs. We analyzed the weight, activation, weight gradient and activation gradient tensors of the feed forward layers, FFN1 and FFN2. Overall, there are 18 x 64 = 1152 shards of each tensor type e.g., FFN1 activation. We present our observations and results for FFN1 activation tensor at data type BFloat16 [25].
4 Observations
Fig. 1 shows the Probability Mass Function (PMF) of FFN1 activation with a symbol size of 8 bits i.e., 256 symbols, averaged over all shards. The probabilities are sorted in decreasing order. Fig. 2 shows the corresponding Cumulative Distribution Function (CDF). The key observation is that the cumulative probability of the top 8 most frequent symbols is about 0.5. The histograms of other tensors, e.g, FFN2 activation, exhibit similar behaviour. This distribution has a Shannon entropy [11] of 6.26 bits and hence an ideal compressibility of $(8-6.26)/8\approx 21.7\%$.
Fig. 3 shows the Huffman [13] code lengths of the different symbols. More frequent symbols have smaller codes and less frequent symbols have longer codes as expected. The compressibility achieved using Huffman codes is 21.3%. For this distribution, the code lengths range from 3 to 10. This makes the Huffman tree very deep and increases the hardware complexity.
Figure 1: Sorted Probability Mass Function (PMF) of FFN1 activation.
Figure 2: Cumulative Distribution Function (CDF) of FFN1 activation.
Figure 3: Huffman code lengths for each symbol.
5 Dual Length Codes
We took a cue from the code lengths assigned by the Huffman algorithm, as shown in Fig. 3. We observed that most symbols have a code length of 9 bits, while a small number of frequently occurring symbols have varying code lengths. This contributes to the optimal compressibility achieved by Huffman codes, but also to its complexity.
We divide the 256 symbols into 2 areas. Identifying the area requires 1 bit. Each area has a different number of symbols. The first area has 8 symbols and hence requires 1 + 3 = 4 bits. The second area has the remaining 248 symbols and hence requires 1 + 8 = 9 bits. This is shown in Fig. 4 and Table 1. The symbol codes are also shown below.
Symbol Code
0      0_000
1      0_001
2      0_010
3      0_011
4      0_100
5      0_101
6      0_110
7      0_111
8      1_0000_1000
9      1_0000_1001
254    1_1111_1110
255    1_1111_1111
Overall, this scheme has codes of 2 different lengths, hence dual length codes. The ‘area code’ i.e., the first bit, encodes the length of the code which is either 4 or 9 bits. In contrast, the Huffman codes shown above have 8 different code lengths. This simplifies the software or hardware implementation. The compressibility achieved using this scheme is 18.6% which is about 2.7% less than Huffman. The compressibility can also be approximated analytically. The top-8 symbols occur with a probability of $\approx$ 0.5 and have a code length of 4 bits. The remaining symbols have a code length of 9 bits, therefore, the expected bits per symbol $0.5\times 4+0.5\times 9=6.5$ bits. Therefore, the compressibility is $(8-6.5)/8=18.75\%$.
Table 1: Dual length coding scheme.
Area Area code #Symbols #Symbol bits Code length Symbol Range
1 0 8 3 4 0-7
2 1 248 8 9 8-255
Figure 4: Code lengths for Huffman codes and dual length codes.
6 Implementation
Fig. 1 shows the sorted PMF of FFN1 activations. Fig. 5 shows the actual probability distribution. The top-8 symbols with the highest probabilities are 61, 62, 63, 64, 189, 190, 191 and 192. In a software or hardware implementation, the encoder and decoder maintain a small Look Up Table (LUT) with codes only for these 8 symbols, as shown below. The remaining symbols are encoded as 1 followed by the binary representation of the symbol in 8 bits, e.g., 35 is encoded as 1_001000111\_00100011. The longer codes are not stored. During encoding, if the symbol exists in the LUT, we use the corresponding 4 bit code, else we use the longer 9 bit code. During decoding, if the leading bit is 0, we read the next 3 bits and then use the LUT for the inverse lookup. If the leading bit is 1, we read the next 8 bits to recover the symbol. Note that multiple LUTs, one for each tensor type e.g., FFN1 activation, FFN1 activation gradient etc., can be obtained apriori as shown in [1].
Symbol Code
61     0_000
62     0_001
63     0_010
64     0_011
189    0_100
190    0_101
191    0_110
192    0_111
Figure 5: Probability Mass Function (PMF) of FFN1 activation.
7 Conclusion
Lossless compression is an effective way to reduce network traffic and improve collective performance. Huffman codes are optimal entropy codes but suffer from slow bit-sequential decoding. This paper introduced Dual Length Codes, a coding scheme for lossless compression of BFloat16 data type in ML applications. The scheme exploits the symbol frequency distributions similar to Huffman codes but trades off some of its compression efficiency for decoding speed and implementation simplicity.
By analyzing the BFloat16 tensors from the Gemma model, we identified that the top-8 most frequent symbols account for approximately 50% of the cumulative probability. These 8 symbols are assigned a short 4 bit code. The remaining 248 symbols are assigned a longer 9 bit code. The codes have a simple encoding and decoding scheme and use a single prefix bit to distinguish between the two code lengths. Unlike Huffman codes which can require deep tree traversals during decoding, this scheme uses a small Look Up Table (LUT) with only 8 entries, significantly reducing the implementation complexity. While a compressibility of 18.6% achieved using these codes is slightly lower than the 21.3% achieved by Huffman codes, Dual Length Codes reduce the latency and complexity associated with Huffman decoding.