### v2

# Quad Length Codes for Lossless Compression of e4m3

Abstract
Training and serving Large Language Models (LLMs) relies heavily on parallelization and collective operations, which are frequently bottlenecked by network bandwidth. Lossless compression using e.g., Huffman codes can alleviate the issue, however, Huffman codes suffer from slow, bit-sequential decoding and high hardware complexity due to deep tree traversals. Universal codes e.g., Exponential-Golomb codes are faster to decode but do not exploit the symbol frequency distributions. To address these limitations, this paper introduces Quad Length Codes, a hybrid approach designed to balance compression efficiency with decoding speed. The coding scheme uses 3 prefix bits to divide the 256 symbols into 8 areas. Each area has a different code length and encodes a different number of symbols. The scheme uses a Look Up Table with 256 entries, significantly simplifying the hardware implementation compared to Huffman trees. The coding scheme can be adapted for different distributions. For the e4m3 data type, the scheme achieves a compressibility of 13.9% in comparison to 15.9% achieved by Huffman codes, but it significantly speeds up the decoding and simplifies the hardware complexity.
1 Background
Training and serving Large Language Models (LLMs) e.g., Gemini [4], Gemma [17, 21], LLaMA [24], GPT [19] require partitioning (sharding) the data (parameters, activations, optimizer state etc.) and parallelizing the computation across multiple accelerators. There are multiple paradigms of parallelism e.g., Data Parallelism, Tensor Parallelism, Pipeline Parallelism, Expert Parallelism and Sequence Parallelism [23, 15, 20, 18, 26, 16, 5]. Different parallelization strategies invoke different collective operations e.g., AllReduce, ReduceScatter, AllGather, AlltoAll [6]. Collective operations are typically bounded by network bandwidth. Lossless compression is an effective way to reduce the network traffic and improve collective performance.
Huffman codes [14] either directly or as part of other algorithms e.g., DEFLATE [8], Zstandard [7], Brotli [3] are commonly used for lossless data compression. They exploit the distribution of symbol frequencies and are optimal entropy codes. However, they are variable length and decoding them requires a binary tree traversal. The decoding is slow and bit sequential i.e., the decode latency is proportional to the number of bits in the encoded input.
Universal codes [25] e.g., Elias Gamma codes [10], Elias Delta codes [9], Elias Omega codes [11], Exponential Golomb codes [13] are also variable length codes. In these codes, the length of the code is part of the code, for example, Elias Gamma codes use a series of leading zeros to indicate the length of the code. Therefore, the decoding is not completely bit sequential and hence faster. However, these codes do not exploit the distribution of symbol frequencies and hence are not optimal.
2 Objective
Our objective is to develop codes such that
  • They exploit the distribution of symbol frequencies, similar to Huffman codes.
  • The decoder is not bit sequential and hence faster, similar to Universal codes.
3 Experimental Setup
We analyzed the Gemma 2B model [17] during Supervised Fine Tuning (SFT). The model has 18 layers and is sharded over 64 TPUs. We analyzed the weight, activation, weight gradient and activation gradient tensors of the feed forward layers, FFN1 and FFN2. Overall, there are 18 x 64 = 1152 shards of each tensor type e.g., FFN1 activation. We present our observations and results for FFN1 and FFN2 activation tensors at the eXmY [1] e4m3 data type where all 256 encodings are finite. The OCP e4m3 format [22] has 2 out of 256 encodings reserved for NaNs, which will have minimal effect on the symbol probabilities and the results. The quantization block size is 32.
4 Observations
Fig. 1 shows the Probability Mass Function (PMF) of FFN1 activation with a symbol size of 8 bits i.e., 256 symbols, averaged over all shards. The probabilities are sorted in decreasing order. The histograms of FFN1 weights, FFN1 weight gradient, FFN2 weights, and FFN2 weight gradient exhibit similar behaviour. This distribution has a Shannon entropy [12] of 6.69 bits and hence an ideal compressibility of $(8-6.69)/8\approx 16.3\%$.
Fig. 2 shows the Huffman [14] code lengths of the different symbols. More frequent symbols have smaller codes and less frequent symbols have longer codes as expected. The compressibility achieved using Huffman codes is 15.9%. For this distribution, the code lengths range from 6 to 18. This makes the Huffman tree very deep and increases the hardware complexity.
Figure 1: Sorted Probability Mass Function (PMF) of FFN1 activation.
Figure 2: Huffman code lengths for each symbol.
5 Quad Length Codes
We took a cue from the code lengths assigned by the Huffman algorithm, as shown in Fig. 2. We observed that the top 37 symbols have a code length of 6 bits, the next 35 symbols have a code length of 7 bits and so on. The other symbols have code lengths from 8 to 18 bits. This contributes to the optimal compressibility achieved by Huffman codes, but also to its complexity.
We divide the 256 symbols into 8 areas. Identifying the area requires 3 bits. Each area has a different number of symbols. The first five areas have 8 symbols each and hence require $3+3=6$ bits, the sixth area has 16 symbols and hence requires $3+4=7$ bits, the seventh area has 32 symbols and hence requires $3+5=8$ bits, and the last area has the remaining 168 symbols and hence requires $3+8=11$ bits. This is shown in Fig. 3 and Table 1.
Overall, this scheme has codes of 4 different lengths, hence quad length codes. The ‘area codes’ i.e., the first three bits, encode the length of the code which can be either 6, 7, 8, or 11 bits. In contrast, the Huffman codes shown above have 13 different code lengths. This simplifies the software or hardware decoder implementation. The compressibility achieved using this scheme is 13.9% which is about 2.0% less than Huffman.
Table 1: Quad length coding scheme.
Area Area code #Symbols #Symbol bits Code length Symbol Range
1 000 8 3 6 0-7
2 001 8 3 6 8-15
3 010 8 3 6 16-23
4 011 8 3 6 24-31
5 100 8 3 6 32-39
6 101 16 4 7 40-55
7 110 32 5 8 56-87
8 111 168 8 11 88-255
Figure 3: Code lengths for Huffman codes and quad length codes.
6 Adaptation
Fig. 4 shows the PMF of FFN2 activation. The probabilities are sorted in decreasing order. The PMF shows that 1 symbol (zero) occurs with a significantly higher frequency than other symbols. This is due to the intervening non-linear activation function. The PMF of FFN1 activation gradient and FFN2 activation gradient are similar. The Shannon entropy of this distribution is 6.11 bits and hence has an ideal compressibility of $(8-6.11)/8=23.6\%$.
Fig. 5 shows the corresponding Huffman code lengths of the different symbols. The code lengths range from 3 to 39 bits. This makes the Huffman tree very deep and increases the hardware complexity. The compressibility achieved using Huffman codes is $23.2\%$.
Using the quad length coding scheme from Table 1 for this distribution results in a compressibility of only $16.7\%$. This is because the frequency distributions are not similar e.g., for this distribution, the most frequent symbol is assigned a code length of 3 bits instead of 6 bits, as in Table 1.
We can adapt the coding scheme to better fit this distribution and is shown in Table 2 and Fig. 6. This coding scheme also has codes of 4 different lengths viz., 4, 6, 8 and 11 bits. The compressibility achieved using this new coding scheme is 19.0%, an improvement of $2.3\%$ from the scheme in Table 1 and about $4\%$ less than Huffman.
Figure 4: Sorted Probability Mass Function (PMF) of FFN2 activation.
Figure 5: Huffman code lengths for each symbol.
Table 2: Quad length coding scheme.
Area Area code #Symbols #Symbol bits Code length Symbol Range
1 000 2 1 4 0-1
2 001 8 3 6 2-9
3 010 8 3 6 10-17
4 011 8 3 6 18-25
5 100 8 3 6 26-33
6 101 32 5 8 34-65
7 110 32 5 8 66-97
8 111 158 8 11 98-255
Figure 6: Code lengths for Huffman codes and quad length codes.
7 Implementation
Fig. 1 and Fig. 4 show the sorted PMF of FFN1 and FFN2 activations. Fig. 7 shows the actual probability distribution of FFN1 activation. The symbols with the highest frequencies are 113, 241, 234, 106 etc. The symbols with the lowest frequencies are 141, 137, 0, 128 etc.
Figure 7: Probability Mass Function (PMF) of FFN1 activation.
In a software or hardware implementation, the encoder maintains a 256 entry Look Up Table (LUT) as shown in Table 3. While creating the table, the input symbols are sorted in the order of decreasing probabilities. They are mapped to symbols 0 through 255 and assigned a code according to the coding scheme of Table 1 or Table 2. After the table is populated, the input symbols i.e., the first column is sorted in ordinal fashion to facilitate indexing into the LUT. The second column is shown only for convenience. During encoding, we simply lookup the code corresponding to the input symbol.
Table 3: Encoder Look Up Table.
Input Symbol Mapped to Symbol Code
113 0 000_000
241 1 000_001
234 2 000_010
233 8 001_000
137 253 111_11111101
0 254 111_11111110
128 255 111_11111111
During decoding, we use the ‘area code’ i.e., the first three bits to determine the length of the code. Using the coding scheme from Table 1, if the area code is $000$ through $100$, we read the next 3 bits. If the area code is $101$, we read the next 4 bits and so on. Then we add the appropriate offset to obtain the encoded symbol. For example, if the area code is $100$ and the next 3 bits are $010$, then the encoded symbol is $32+2=34$. The decoder maintains a 256 entry LUT to convert the encoded symbol to the output symbol as shown in Table 4. Note that multiple LUTs, one for each tensor type e.g., FFN1 activation, FFN1 activation gradient etc., can be obtained apriori as shown in [2].
Table 4: Decoder Look Up Table.
Encoded Symbol Output Symbol
0 113
1 241
2 234
8 233
253 137
254 0
255 128
8 Conclusion & Future Work
Lossless compression is an effective way to reduce network traffic and improve collective performance. Huffman codes are optimal entropy codes but suffer from slow bit-sequential decoding. This paper introduced Quad Length Codes, a coding scheme for lossless compression of e4m3 data type in ML applications. The scheme exploits the symbol frequency distributions similar to Huffman codes but trades off some of its compression efficiency for decoding speed and implementation simplicity.
The coding scheme uses 3 prefix bits to divide the 256 symbols into 8 areas. Each area has a different number of symbols and hence has a different code length. The scheme uses a Look Up Table with 256 entries, significantly reducing the implementation complexity compared to Huffman trees. While a compressibility of 13.9% achieved using these codes is slightly lower than the 15.9% achieved by Huffman codes, Quad Length Codes reduce the latency and complexity associated with Huffman decoding.
Our coding schemes were obtained empirically. It is possible to tweak the number of areas, the number of symbols in each area, and the number of unique code lengths to achieve a better compression ratio for the above and other distributions. In future, we want to develop a mathematical formulation of the problem.